Can someone solve #6, that’s the one I don’t understand
Phys 11A HW1 to grade Find the components of the three vectors as shown. a-8m, b-6m, c-5m. Find a +b-c both, graphically and analytically 1. 30° 40° 20 2. Findi j+ k in unit-vector notations. Is it a unit-vector? Justify your answer 3. Vector a is 3m long and is 60° above x-axis in the first quadrant. Vector b is 5m long and is 50° below the x-axis in the fourth quadrant. Find aja + b, ba-b, c) b-a. Provide answers to a)-c) in both, unit-vector notations and in terms of magnitude and direction. 4. For two vectors mentioned in the problem 3, find a b and axb (in terms of magnitude and direction). 5. Find the angle between two vectors а-6i-4) and b 6. For three vectors mentioned in problem 1, find ä (bxč). -10-31
Answer #1

Problem -6

Given is:-

magnitude of vectors a=8m,b=6m,c=5m

Now,

the components of the vectors according to the diagram shown is

a 8cos(40) (6.13i)m y 8sin(40) (5.14j)m thus i-(6.13i 5.147)m

similarly

6-6cos(30) (-5.2)m 6sin(30 (3)m hus

and

5sin(20-1.71i)m Cy--5cos(20) (-4.7j)m COS hus (-1.71i-4.7)m

Now,

c= (-52i + 3j) × (-1.71 i-4.77)

which gives us

b × c= (29.571)

now

ī.(b × C) = (6.132 + 5.14j ) . (29.57k)

which gives us

oxed{vec a . (vec b imes vec c) = 0}